OpenStudy (anonymous):

The graph below shows the price, y, in dollars, of different amounts of peanuts, x, in pounds: A graph titled Peanut Prices shows Number of Pounds on x axis and Price in dollars on y axis. The scale on the x axis shows numbers from 0 to 12 at increments of 2, and the scale on the y axis shows numbers from 0 to 36 at increments of 6. A straight line joins the ordered pairs 0, 0 and 12, 36. Which equation best represents the relationship between x and y? y = x + 6 y = 3x y = 6x y = x + 3

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

@HelpOfTheGods @help?!?!?! @teddyiswhathecallsme. @sammixboo

OpenStudy (anonymous):

@JackofallTradez

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (0,0)...x1 = 0 and y1 = 0 (12,36)...x2 = 12 and y2 = 36 now we sub slope(m) = (36 - 0) / (12 - 0) slope(m) = 36/12 = 3 y = mx + b slope(m) = 3 use either points (0,0)...x = 0 and y = 0 now sub 0 = 3(0) + b 0 = b so your equation is : y = 3x + 0....or just y = 3x

OpenStudy (texaschic101):

questions ?

OpenStudy (anonymous):

THANKS U SOOOOOOOOOOOOOOOO MUCHHHHHHHHHHHHH @texaschic101

OpenStudy (texaschic101):

no problem :)